# An anomaly-free leptophilic $U(1)^{\prime}_{\ell}$ completion of the Standard Model

### v1

@@ Abstract
The model thus provides a single, internally consistent benchmark that unifies gauge, scalar, Yukawa, neutrino, and dark-matter sectors, and is directly suited for global data recasts and targeted searches at future lepton and hadron colliders.
@@ I Introduction
Neutrino oscillations [1] have established that individual lepton flavors are not conserved, implying the existence of physics beyond the Standard Model (SM). A minimal way to address this is to enlarge the gauge sector with an additional Abelian symmetry that acts on leptons. In this spirit, we consider a general leptophilic $U(1)^{\prime}_{\ell}$ extension in which the three SM lepton families carry arbitrary charges $(q_{e},q_{\mu},q_{\tau})$ while quarks are neutral. Notable special cases include the universal choice $q_{e}=q_{\mu}=q_{\tau}$, as well as the well-known $L_{\mu}-L_{\tau}$ scenario [2]. The associated gauge boson, denoted $Z^{\prime}_{\ell}$ (or $Z_{\ell}$ henceforth), couples to leptons at tree level and acquires quark couplings only through kinetic mixing.
Extensions of the SM by extra $U(1)$ factors remain among the simplest and most predictive portals for new interactions [3]. A fully consistent realization, however, requires a definite ultraviolet (UV) completion in which gauge, mixed, and gravitational anomalies are cancelled and the symmetry is spontaneously broken. In this work we formulate such a completion for a generic leptophilic $U(1)^{\prime}_{\ell}$, using one vectorlike lepton family and SM-singlet scalars whose vacuum expectation values generate masses for both $Z_{\ell}$ and the heavy leptons. We work in a left-handed Weyl basis and derive compact anomaly-cancellation relations applicable to both universal and non-universal charge assignments.
Kinetic mixing between $U(1)_{Y}$ and $U(1)^{\prime}_{\ell}$ [4] is automatically induced at the loop level unless $\mathrm{Tr}(YQ^{\prime})=0$, a condition naturally satisfied for several anomaly-free solutions. We present analytic $Z_{\ell}$ couplings and widths, including the effects of kinetic mixing, and discuss useful limits such as the contact-interaction regime $M_{Z_{\ell}}^{2}\gg s$ [5]. The setup accommodates type-I seesaw realizations for neutrino masses and offers minimal dark-matter candidates in the form of stable singlet fermions annihilating through $Z_{\ell}$ exchange [6]. It therefore ties collider, precision, and cosmological probes to a single gauge parameter space.
The goal of this paper is analytic and model-building oriented: to provide a closed-form benchmark that links gauge, scalar, and Yukawa sectors, and that enables transparent phenomenological mappings. We present the field content, charge assignments, anomaly checks, symmetry breaking, mass spectra, kinetic-mixing structure, and resulting parametric dependences for collider, precision, and dark-sector observables. A full global analysis (including likelihood fits and public code) is left to future work.
@@ II Gauge structure and anomaly cancellation
Conventions. We use $Q=T_{3}+Y$, with $Y_{\rm PDG}=2Y$. The covariant derivative is $D_{\mu}=\partial_{\mu}-igW_{\mu}^{a}T^{a}-ig_{Y}YB_{\mu}-ig_{\ell}Q^{\prime}B^{\prime}_{\mu}$, and kinetic mixing $\varepsilon$ is introduced in Sec. V.
@@
where $U(1)^{\prime}_{\ell}$ acts on leptons but not on quarks. We adopt the standard hypercharge normalization $Q=T_{3}+Y$ so that $Y=Y_{\rm PDG}/2$.
Before turning to the most general flavor-dependent case, it is illuminating to begin with a universal charge assignment and examine the ensuing anomaly structure.
@@ II.1 Universal assignment and anomaly constraints
while allowing the Higgs to remain uncharged. Unlike the special choice $L_{\mu}-L_{\tau}$ [7, 8], this universal assignment generates mixed anomalies with $SU(2)_{L}$ and $U(1)_{Y}$. A consistent $U(1)^{\prime}$ gauge theory requires these anomalies to vanish, otherwise gauge invariance would be broken at the quantum level.
@@
The mixed anomalies must therefore be cancelled by new fermions. Guided by the principle of minimality—namely keeping the exotic content to a single vectorlike family while retaining anomaly freedom—we introduce one vectorlike lepton doublet $(L^{\prime}_{L},L^{\prime}_{R})$ and one vectorlike charged singlet $(E^{\prime}_{L},E^{\prime}_{R})$, vectorlike under the SM but chiral under $U(1)^{\prime}_{\ell}$. [^1] See Appendix A for charge conventions and detailed derivations of the anomaly conditions.
@@ II.2 Complete anomaly cancellation for general $(q_{e},q_{\mu},q_{\tau})$
the anomaly-free assignments for the exotics are:
@@
where $\Delta_{L,E}\equiv Q^{\prime}(L^{\prime}_{L,E})-Q^{\prime}(L^{\prime}_{R,E})$ and $s_{L,E}\equiv Q^{\prime}(L^{\prime}_{L,E})+Q^{\prime}(L^{\prime}_{R,E})$. These cancel $SU(2)_{L}]^{2}U(1)^{\prime}_{\ell}$, $U(1)_{Y}]^{2}U(1)^{\prime}_{\ell}$, and gravitational anomalies, while the cubic condition fixes $s_{L}$:
@@
This construction illustrates that anomaly-free leptophilic $U(1)^{\prime}$ models are highly constrained yet flexible enough to encode both universal and flavor-dependent scenarios. Further details and example charge tables are given in Appendix A.
@@ IV.2 Two-singlet generalization
If two scalars $\Phi_{1,2}$ with charges $q_{1,2}$ acquire vevs $v_{1,2}$, gauge invariance requires
@@
so that $L^{\prime}$ masses arise from $\Phi_{1}$ while $E^{\prime}$ and $N^{\prime}$ receive masses from $\Phi_{2}^{\dagger}$. The resulting spectrum is
$m_{L^{\prime}}=\frac{y_{L}v_{1}}{\sqrt{2}},\qquad m_{E^{\prime}}=\frac{y_{E}v_{2}}{\sqrt{2}},\qquad m_{N^{\prime}}=\frac{y_{N}v_{2}}{\sqrt{2}},$ (11)
allowing for non-degenerate exotic-lepton masses while preserving anomaly cancellation.
@@ V Kinetic mixing and precision constraints
and $\varepsilon$ remains threshold-suppressed and naturally small. Throughout we adopt the convention $Q=T_{3}+Y$.
@@ VI.2 Phenomenology and cosmology
For $v_{S}\!\sim\!{\rm TeV}$, the heavy neutrinos may be accessible at colliders through displaced-vertex or multilepton signatures [16, 17, 18]. For larger $v_{S}$, the seesaw scale decouples but still permits thermal leptogenesis [19, 20]. Because $\nu^{c}$ carries $Y=0$, sterile states do not contribute to $\mathrm{Tr}(YQ^{\prime})$ and thus do not upset the kinetic-mixing structure of Sec. V. The neutrino sector therefore provides a second, independent probe of the $U(1)^{\prime}_{\ell}$ breaking scale, linking collider, precision, and cosmological signatures.

### v2

@@ Abstract
The model thus provides a single, internally consistent benchmark that unifies gauge, scalar, Yukawa, neutrino, and dark-matter sectors, and is directly suited for global data recasts and targeted searches at current and future lepton and hadron colliders
@@ I Introduction
Neutrino oscillations [1] have established that individual lepton flavors are not conserved, implying physics beyond the Standard Model (SM). A minimal way to address this is to gauge an additional Abelian symmetry that acts on leptons. In this spirit, we consider a general leptophilic $U(1)^{\prime}_{\ell}$ extension in which the three SM lepton families carry arbitrary charges $(q_{e},q_{\mu},q_{\tau})$ while quarks are neutral. Prominent special cases include the universal assignment $q_{e}=q_{\mu}=q_{\tau}$ and the well-known $L_{\mu}-L_{\tau}$ scenario [2]. The associated gauge boson, denoted $Z^{\prime}_{\ell}$ (or $Z_{\ell}$ henceforth), couples to leptons at tree level, while any quark coupling arises only via kinetic mixing.
Extensions of the SM by extra $U(1)$ factors remain among the simplest and most predictive portals for new interactions [3]. A fully consistent framework, however, requires a UV-complete realization in which gauge, mixed, and gravitational anomalies are canceled and the symmetry is spontaneously broken. Here we construct such a realization for a generic leptophilic $U(1)^{\prime}_{\ell}$ by introducing one vectorlike lepton family and SM-singlet scalars whose vacuum expectation values generate masses for both $Z_{\ell}$ and the heavy leptons. We work in a left-handed Weyl basis and derive compact anomaly-cancellation relations applicable to both universal and nonuniversal charge assignments.
Kinetic mixing between $U(1)_{Y}$ and $U(1)^{\prime}_{\ell}$ [4] is radiatively induced unless $\mathrm{Tr}(YQ^{\prime})=0$, a condition naturally satisfied for several anomaly-free solutions. We present analytic $Z_{\ell}$ couplings and widths, including kinetic-mixing effects, and discuss useful limits such as the contact-interaction regime $M_{Z_{\ell}}^{2}\gg s$ [5]. The setup accommodates type-I seesaw realizations of neutrino masses and admits minimal dark-matter candidates in the form of stable singlet fermions annihilating through $Z_{\ell}$ exchange [6], thereby tying collider, precision, and cosmological probes to a single gauge-parameter space.
The goal of this paper is analytic and model-building–oriented: to provide a closed-form benchmark that links gauge, scalar, and Yukawa sectors and enables transparent phenomenological mappings. We present the field content, charge assignments, anomaly checks, symmetry breaking, mass spectra, kinetic-mixing structure, and resulting parametric dependences for collider, precision, and dark-sector observables. A full global analysis (including likelihood fits and public code) is left to future work.
Models gauging total lepton number have been discussed in various contexts [7, 8, 9, 10, 11, 12]. The present work differs by providing a fully anomaly–free realization with a single vectorlike family, connecting the $U(1)^{\prime}_{\ell}$ breaking scale to neutrino masses and dark matter, and by embedding all phenomenological predictions in a compact $(M_{Z_{\ell}},g_{\rm eff})$ parameter space.
@@ II Gauge structure and anomaly cancellation
Conventions. We use $Q=T_{3}+Y$ with $Y_{\rm PDG}=2Y$. The covariant derivative is $D_{\mu}=\partial_{\mu}-igW_{\mu}^{a}T^{a}-ig_{Y}YB_{\mu}-ig_{\ell}Q^{\prime}B^{\prime}_{\mu},$ and kinetic mixing $\varepsilon$ is introduced in Sec. V.
@@
where $U(1)^{\prime}_{\ell}$ acts on leptons but not on quarks. We adopt the standard hypercharge normalization $Q=T_{3}+Y$, implying $Y=Y_{\rm PDG}/2$.
Before turning to the most general flavor-dependent case, it is instructive to begin with a universal charge assignment and examine the resulting anomaly pattern.
@@ II.1 Universal assignment and anomaly constraints
while keeping the Higgs uncharged. Unlike the special choice $L_{\mu}-L_{\tau}$ [13, 14], this universal assignment induces mixed anomalies with $SU(2)_{L}$ and $U(1)_{Y}$. A consistent $U(1)^{\prime}$ gauge theory requires these anomalies to vanish; otherwise gauge invariance would be violated at the quantum level.
@@
The mixed anomalies must therefore be canceled by new fermions. Guided by the principle of minimality—keeping the exotic content to a single vectorlike family while retaining anomaly freedom—we introduce one vectorlike lepton doublet $(L^{\prime}_{L},L^{\prime}_{R})$ and one vectorlike charged singlet $(E^{\prime}_{L},E^{\prime}_{R})$, vectorlike under the SM but chiral under $U(1)^{\prime}_{\ell}$. [^1] See Appendix A for charge conventions and explicit derivations of the anomaly conditions.
@@ II.2 Complete anomaly cancellation for general $(q_{e},q_{\mu},q_{\tau})$
the anomaly-free assignments for the exotics are
@@
where $\Delta_{L,E}\equiv Q^{\prime}(L^{\prime}_{L,E})-Q^{\prime}(L^{\prime}_{R,E})$ and $s_{L,E}\equiv Q^{\prime}(L^{\prime}_{L,E})+Q^{\prime}(L^{\prime}_{R,E})$. These relations cancel $SU(2)_{L}]^{2}U(1)^{\prime}_{\ell}$, $U(1)_{Y}]^{2}U(1)^{\prime}_{\ell}$, and gravitational anomalies, while the cubic condition fixes $s_{L}$:
@@
This construction shows that anomaly-free leptophilic $U(1)^{\prime}$ models are highly constrained yet sufficiently flexible to realize both universal and flavor-dependent scenarios. Together with the scalar and Yukawa sectors discussed in the following sections, it completes the anomaly-free backbone promised in Sec. I, establishing the theoretical closure that will underpin the phenomenological analysis to come.
Explicit benchmark example.
For illustration, consider the nonuniversal choice
$(q_{e},q_{\mu},q_{\tau})=(1,1,-2),$
for which $\Sigma q=0$ and $\Sigma q^{3}=1+1-8=-6$. In this case all gauge and gravitational anomalies cancel within the SM (augmented by $\nu^{c}_{i}$) and no exotic fermions are required, as anticipated by the corollary below Eq. (12). By contrast, for a nonvanishing trace, e.g.
$(q_{e},q_{\mu},q_{\tau})=(1,0,0),$
one finds $\Sigma q=1$ and the anomaly conditions fix the chiral differences $\Delta_{L,E}$ and the sums $s_{L,E}$ as in Eq. (11), together with the cubic constraint on $s_{L}$. The corresponding Yukawa structure follows from the same charge bookkeeping: for each lepton family the admissible $\Phi$-Yukawa couplings are obtained by solving $Q^{\prime}(F_{L})+Q^{\prime}(F_{R})+Q^{\prime}(\Phi)=0$, so that mass terms for the exotic states are again directly tied to the $U(1)^{\prime}_{\ell}$ breaking pattern. This makes explicit how the general framework smoothly interpolates between universal and family-dependent leptophilic scenarios.
@@ IV.2 Two-singlet generalization
If two scalars $\Phi_{1,2}$ with charges $q_{1,2}$ acquire vevs $v_{1,2}$, gauge invariance in the universal benchmark requires
@@
so that $q_{1}=q_{2}$ in this specific realization. In other words, both $\Phi_{1}$ and $\Phi_{2}$ are allowed to couple to $(L^{\prime},E^{\prime},N^{\prime})$ at the renormalizable level. In our minimal texture we choose a basis in which $L^{\prime}$ masses arise dominantly from $\Phi_{1}$ while $E^{\prime}$ and $N^{\prime}$ receive masses from $\Phi_{2}^{\dagger}$,
$\mathcal{L}_{Y}\supset y_{L}\,\Phi_{1}\,L^{\prime}L^{\prime c}+y_{E}\,\Phi_{2}^{\dagger}E^{\prime}E^{\prime c}+y_{N}\,\Phi_{2}^{\dagger}N^{\prime}N^{\prime c}+{\rm h.c.},$ (11)
so that
$m_{L^{\prime}}=\frac{y_{L}v_{1}}{\sqrt{2}},\qquad m_{E^{\prime}}=\frac{y_{E}v_{2}}{\sqrt{2}},\qquad m_{N^{\prime}}=\frac{y_{N}v_{2}}{\sqrt{2}}.$ (12)
Additional mixed couplings (e.g. $\Phi_{1}^{\dagger}E^{\prime}E^{\prime c}$ or $\Phi_{2}L^{\prime}L^{\prime c}$) are technically natural to set to zero: they are absent at tree level and only regenerated radiatively at higher order, and including them does not affect anomaly cancellation. We therefore retain the simpler texture above as our benchmark without loss of generality.
@@ IV.3 Mixing structure and flavor protection
Theoretical necessity.
The additional vectorlike leptons introduced above are not an ad hoc embellishment but a minimal and unavoidable ingredient of the construction: they are required to ensure full gauge and gravitational anomaly cancellation in a leptophilic $U(1)^{\prime}_{\ell}$ extension. While such states have not been observed experimentally, their presence is a generic prediction of anomaly-free frameworks of this type and can be probed through multilepton and (potentially) long-lived signatures at colliders. This theoretical necessity underpins the phenomenological study in Sec. VIII, where we map the $(M_{Z_{\ell}},g_{\rm eff})$ plane to the exotic-lepton spectrum and search channels.
@@ IV.4 Couplings and collider outlook
Together with the anomaly-free gauge and scalar sectors discussed above, the Yukawa structure completes the renormalizable foundation of the model, linking the symmetry-breaking scale to the observable lepton spectrum and collider signatures.
@@ V Kinetic mixing and precision constraints
and $\varepsilon$ remains threshold-suppressed and naturally small. [^3] For vectorlike pairs with masses $m_{1}\neq m_{2}$, finite threshold pieces scale as $\Delta\varepsilon\sim(g_{Y}g_{\ell}/16\pi^{2})\,(Y\,\Delta Q^{\prime})\,\ln(m_{1}^{2}/m_{2}^{2})$, yielding $\varepsilon\sim 10^{-4}\text{--}10^{-3}$ for ${\cal O}(1)$ splittings. Throughout we adopt the convention $Q=T_{3}+Y$.
@@ A Majorana mass for the right-handed neutrinos can be generated by a singlet $S$ with
Since $S$ is charged under $U(1)^{\prime}_{\ell}$, its vev also contributes to the $Z_{\ell}$ mass. In the presence of all singlet vevs, the gauge boson mass is
$M_{Z_{\ell}}^{2}=g_{\ell}^{2}\big(q_{\Phi}^{2}v_{\Phi}^{2}+q_{S}^{2}v_{S}^{2}+\cdots\big),$ (21)
which reduces to Eq. (7) when $v_{S}\ll v_{\Phi}$ or when the $S$ contribution is absorbed into an effective $v_{\Phi}$. In our phenomenological benchmarks we are effectively working in this limit, so that $M_{Z_{\ell}}$ can be parameterized in terms of a single breaking scale while the seesaw scale is controlled by $v_{S}$ and $y_{N}$.
@@ VI.2 Phenomenology and cosmology
For $v_{S}\!\sim\!{\rm TeV}$, the heavy neutrinos may be accessible at colliders through displaced-vertex or multilepton signatures [22, 23, 24]. For larger $v_{S}$, the seesaw scale decouples but still permits thermal leptogenesis [25, 26]. Because $\nu^{c}$ carries $Y=0$, sterile states do not contribute to $\mathrm{Tr}(YQ^{\prime})$ and thus do not upset the kinetic-mixing structure of Sec. V. The neutrino sector therefore provides a second, independent probe of the $U(1)^{\prime}_{\ell}$ breaking scale, linking collider, precision, and cosmological signatures. Taken together, the anomaly-free gauge completion, scalar breaking, and neutrino-mass generation establish a coherent leptophilic framework linking particle masses to the same $U(1)^{\prime}_{\ell}$ origin.
@@ VII.3 Direct and indirect searches
Benchmark point and scalar portal.
To illustrate viability, consider a representative fermionic benchmark
$M_{Z_{\ell}}=1~\text{TeV},\quad|g_{\rm eff}|=0.3,\quad g_{\chi}=0.5,\quad m_{\chi}=400~\text{GeV}.$
Away from the ss–channel resonance, Eq. (23) yields an annihilation cross section of order $\langle\sigma v\rangle\sim 10^{-26}\,\text{cm}^{3}/\text{s}$, which is compatible with the observed relic density while satisfying collider and indirect limits for leptonic final states. This explicitly demonstrates that large portions of the parameter space shown in Fig. 1 remain viable once dark–matter constraints are correctly restricted to leptophilic scenarios.
For scalar dark matter, Higgs–portal interactions of the form $\lambda_{XH}\,|X|^{2}|H|^{2}$ are tightly constrained by direct detection. In our benchmarks we set $\lambda_{XH}=0$ at tree level, so that annihilation proceeds dominantly via $Z_{\ell}$ exchange as in Eq. (24), and radiatively induced portal couplings remain well below current experimental limits. This choice is technically natural and consistent with the leptophilic focus of the model.
@@ VIII Phenomenology
Related work.
Complementary phenomenological studies of leptophilic $Z^{\prime}$ scenarios have been performed in global analyses [27], where collider and flavour data were combined to constrain the effective couplings. In contrast, the present work provides a fully anomaly-free and ultraviolet-complete realization of a leptophilic $U(1)^{\prime}_{\ell}$ symmetry, linking the gauge, scalar, and Yukawa sectors in a single predictive framework that can naturally reproduce and extend such phenomenological bounds.
@@ Appendix B Extended scalar sector and mixing
CP–odd sector.
Writing $\Phi_{i}=\tfrac{1}{\sqrt{2}}(v_{i}+\phi_{i}+ia_{i})$, the CP–odd fields $(a_{1},a_{2})$ span one would–be Goldstone boson $G^{\prime}$ and one physical pseudoscalar $A$. In the $(a_{1},a_{2})$ basis, the quadratic terms from $V_{\rm soft}$ give
$\mathcal{M}^{2}_{\rm odd}=\mu_{12}^{2}\begin{pmatrix}\displaystyle\frac{v_{2}}{v_{1}}&-1\\ -1&\displaystyle\frac{v_{1}}{v_{2}}\end{pmatrix}$. (B.11)
The eigenstates are
$G^{\prime}$ $\propto v_{1}a_{1}+v_{2}a_{2},$ (B.12)
$A$ $\propto-v_{2}a_{1}+v_{1}a_{2},$ (B.13)
with
$m_{A}^{2}=\mu_{12}^{2}\left(\frac{v_{1}}{v_{2}}+\frac{v_{2}}{v_{1}}\right),\qquad m_{G^{\prime}}^{2}=0.$ (B.14)
In the limit $\mu_{12}^{2}\to 0$, the CP–odd scalar $A$ becomes massless, signaling the restoration of a global $U(1)$ in the singlet sector. Conversely, a nonzero $\mu_{12}^{2}$ both lifts $A$ and ensures that only one Goldstone boson is eaten by $Z_{\ell}$.